# Towards Practical Benchmarking of Data Cleaning Techniques: On Generating Authentic Errors via Large Language Models

### v2

@@ 7.3. Error Repairing
Error repairing aims to restore data integrity by correcting erroneous values. Traditional methods include statistical reasoning (e.g., ERACER (mayfield2010eracer), SCARE (yakout2013don)) and constraint-based approaches enforcing FDs and CFDs (e.g., NADEEF (ebaid2013nadeef), Holistic (chu2013holistic)). Clustering-based strategies leverage data similarity for correction (e.g., DORC (song2015turn), DISC (song2021saving)).
Recent ML-based methods integrate probabilistic reasoning and external data sources (e.g., HoloClean (rekatsinas2017holoclean), BoostClean (krishnan2017boostclean)) to refine repairs. DL-based approaches enhance repair by learning data dependencies (e.g., DeepClean (zhang2018deepclean)) and leveraging large-scale contextual embeddings (e.g., LLMClean (biester2024llmclean)). While these methods improve repair accuracy, they remain computationally expensive, requiring further optimization (tong2016online).

### v3

@@ 7.3. Error Repairing
Error repairing aims to restore data integrity by correcting erroneous values. Traditional methods include statistical reasoning (e.g., ERACER (Mayfield et al., 2010), SCARE (Yakout et al., 2013)) and constraint-based approaches enforcing FDs and CFDs (e.g., NADEEF (Ebaid et al., 2013), Holistic (Chu et al., 2013b)). Clustering-based strategies leverage data similarity for correction (e.g., DORC (Song et al., 2015a), DISC (Song et al., 2021)).
Recent ML-based methods integrate probabilistic reasoning and external data sources (e.g., HoloClean (Rekatsinas et al., 2017), BoostClean (Krishnan et al., 2017)) to refine repairs. DL-based approaches enhance repair by learning data dependencies (e.g., DeepClean (Zhang et al., 2018)) and leveraging large-scale contextual embeddings (e.g., LLMClean (Biester et al., 2024)). While these methods improve repair accuracy, they remain computationally expensive, requiring further optimization (Tong et al., 2016).